### v1

# OPENXRD: A Comprehensive Benchmark and Enhancement Framework for LLM/MLLM XRD Question Answering

[…]
tcb@breakable
[…]
OPENXRD: A Comprehensive Benchmark and Enhancement Framework for LLM/MLLM XRD Question Answering
Ali Vosoughi,a Ayoub Shahnazari,b Yufeng Xi,c Zeliang Zhang,a Griffin Hess,b Chenliang Xua and Niaz Abdolrahimb,c,d
This work presents OPENXRD, an open-book pipeline designed for crystallography question answering, which integrates textual prompts with concise supporting content generated by GPT-4.5. Instead of using scanned textbooks, which may lead to copyright issues, OPENXRD generates compact, domain-specific references that help smaller models understand key concepts in X-ray diffraction (XRD). We evaluate OPENXRD on a well-defined set of 217 expert-level XRD questions by comparing different vision-language models, including GPT-4 and LLaVA-based frameworks such as Mistral, LLaMA, and QWEN, under both closed-book (without supporting material) and open-book (with supporting material) conditions. Our experimental results show significant accuracy improvements in models that use the GPT-4.5-generated summaries, particularly those with limited prior training in crystallography. OPENXRD uses knowledge from larger models to fill knowledge gaps in crystallography and shows that AI-generated texts can help smaller models reason more effectively in scientific tasks. While the current version of OPENXRD focuses on text-based inputs, we also explore future extensions such as adding real crystal diagrams or diffraction patterns to improve interpretation in specialized materials-science contexts. Overall, OPENXRD shows that specialized open-book systems can be useful in materials science and provides a foundation for broader natural language processing (NLP) tools in critical scientific fields.
footnotetext:  These authors contributed equally to this work.footnotetext: a Department of Computer Science, University of Rochester, Rochester, New York 14627, USA.footnotetext: b Department of Mechanical Engineering, University of Rochester, Rochester, 14627, NY, USA. footnotetext: c Material Science Program, University of Rochester, Rochester, 14627, NY, USA. footnotetext: d Staff Scientist, Laboratory for Laser Energetics (LLE), University of Rochester, Rochester, 14627, NY, USA. footnotetext: † The codes and benchmarking dataset can be accessed at https://github.com/niaz60/OpenXRD. Project webpage: https://niaz60.github.io/OpenXRD/.
1 Introduction
Crystallography is the scientific discipline concerned with determining the arrangement of atoms and molecules in crystalline solids 1, 2, 3. This structural understanding is crucial for elucidating material properties, such as symmetry, geometry, and physical characteristics 4, 5. This understanding is crucial for progressing materials science, especially in areas like metallurgy, pharmaceuticals, and semiconductor technology 6, 7, 8, 9, 10. Central to crystallography is X-ray diffraction (XRD), a key experimental technique that enables researchers to uncover detailed information about crystal structures by examining how X-rays interact with crystalline lattices 11, 12, 13.
The foundational principle of XRD is Bragg’s Law, discovered by Lawrence Bragg in 1912 14, 15. Bragg’s law describes how the angle of a diffracted X-ray beam depends on the wavelength of the X-rays and the spacing between atoms and molecules within the material 16, 17. Using XRD, researchers analyze crystal structures to determine important parameters such as atomic arrangements, phase composition, unit cell size, grain size, crystallinity, strain, and lattice defects 18, 19, 20, 21, 22, 23.
Advances in the analysis of large XRD data have demonstrated that conventional deep learning approaches, such as Convolutional Neural Networks (CNNs) and Graph Neural Networks (GNNs), are powerful tools, particularly for identifying or predicting crystal parameters space group labels 24, 25, 26, lattice constants 27, 28, or phase compositions 29, 30. These models achieve remarkable numerical accuracy, demonstrating their efficacy for crystal classification and structural analysis tasks.
However, despite their success in quantitative predictions, these deep-learning techniques are limited in providing interpretative and explanatory insights into the underlying physics or chemistry of XRD data 31, 32, 33. This lack of interpretability restricts their ability to offer meaningful explanations about the material properties or structural characteristics encoded within the diffraction data. This lack of interpretability restricts the broader application of deep-learning approaches in crystallographic research, where understanding underlying phenomena is essential 34, 35, 36.
In contrast, recent progress in Natural Language Processing (NLP) driven by Large Language Models (LLMs), such as GPT-based architectures, has transformed various computational domains 37, 38, 39. LLMs demonstrate exceptional capabilities in open-ended question answering 40, 41, multi-step reasoning 42, 43, domain-specific question answering 44, 45, 46, predicting material properties 47, 48, 49, and extracting information from complex datasets 50, 51, 52. These powerful language models can potentially bridge the interpretability gap faced by traditional deep-learning models when applied to scientific fields, including materials science and crystallography 53, 54, 55.
For instance, Antunes et al. introduced CrystaLLM, an autoregressive LLM trained on millions of Crystallographic Information Files (CIFs), which encode detailed crystallographic data, including atomic coordinates, symmetry operations, lattice parameters, and space group information, to generate plausible crystal structures from given chemical compositions, demonstrating the potential of text-based crystal structure generation 56. In addition, Johansen et al. introduced deCIFer, an autoregressive language model on powder XRD data to produce complete crystal structures in CIF format, attaining high match rates between generated structures and experimental diffraction profiles 57.
Further extending these capabilities, Choudhary introduced DiffractGPT, a generative pretrained transformer that predicts atomic structures directly from XRD patterns, particularly when chemical information is provided 58. AtomGPT, another transformer-based model by Choudhary, effectively predicts material properties such as formation energies, electronic bandgaps, and superconducting transition temperatures with accuracy comparable to GNNs 59. Beyond generative tasks, LLMs have also been employed for scientific question-answering and knowledge retrieval in the materials domain. For example, the LLaMP framework combines an LLM with a materials database to fetch and reason over crystallographic data, reducing hallucinations and improving factual accuracy in materials science QA 51.
In this study, we investigate whether providing domain-specific context generated by a stronger model, such as supporting textual material, can significantly enhance the performance of weaker models or alternative configurations of the same model on specialized XRD-related questions. To rigorously test this hypothesis, we constructed a carefully selected dataset consisting of 217 multiple-choice XRD questions, each reviewed and approved by a domain expert at the Ph.D. level. Each question has only one correct answer among four provided choices, covering a range of topics from fundamental principles to complex scenarios, including issues such as basic structural geometry, unit cell dimensions, and coordination environments, to more complex concepts, including fundamental equations and symmetry analysis. Initially, we evaluate the performance of various Large LLMs, including GPT-4.5, GPT-4, O1, O3, and so forth, under closed-book conditions, where models rely solely on their pretrained internal knowledge without external assistance. Our initial results show that GPT-4.5 clearly outperforms other tested models. Based on this outcome, we employ GPT-4.5 to generate approximately one-page textual summaries. These summaries intentionally avoid explicitly stating correct answers yet offer adequate context and guidance to facilitate correct reasoning. We then reassess the performance of the other models under open-book conditions to quantify the improvement when provided with these contextual summaries, aiming to quantify the improvement enabled by this supplementary textual material.
2 Methods
2.1 Dataset selection and model processing
To evaluate language model performance in crystallography, we curated a domain-specific multiple-choice question-answering dataset, comprising 217 items focused exclusively on X-ray diffraction and crystallography. To construct this dataset, we partially drew on content from Cullity’s Introduction to X-ray Diffraction 11. Each question in the dataset includes a concise prompt, a set of three to four answer options, a single correct answer, a brief explanation providing the rationale, and a corresponding subtask label. These subtasks span a range of crystallographic domains, including crystal structure fundamentals, diffraction geometry, unit cell analysis, and related topics. As illustrated in Figure 1, the distribution of subtask labels highlights the diversity and complexity of the topics covered in our dataset, ranging from basic definitions to advanced structural phenomena. All questions were reviewed and validated by a domain expert to ensure scientific correctness and comprehensive coverage of essential XRD subtasks.
Fig.  1: A word cloud of the subtask labels from our human-curated crystallography dataset. Larger words indicate subtasks with a higher number of questions, illustrating the breadth of topics (e.g., diffraction fundamentals, lattice geometry, advanced structure analysis). This distribution underscores the diversity and domain complexity captured in our QA benchmark, spanning both basic definitions (e.g., counting crystal systems) and intricate reflections (e.g., twin boundaries, space-group anomalies).
This annotated dataset serves three primary purposes. First, it provides a benchmarking framework for testing the capability of LLMs to comprehend and apply specialized crystallographic concepts. Second, it enables fine-grained subtask-level analysis, distinguishing model performance across different categories, ranging from fundamental definitions to complex conditions. Third, it supports a comparative study between closed-book and open-book evaluation settings, allowing us to assess the impact of supplementary domain-specific context on model accuracy.
AI-Generated Material The atomic scattering factor measures how effectively an atom scatters X-rays, and it depends on the atom’s electron density and the scattering angle. Expert-Reviewed Material The atomic scattering factor quantifies how effectively an atom scatters X-rays, depending primarily on its electron density distribution and the scattering angle. At low scattering angles, the atomic scattering factor approaches the total electron count of the atom, but as the angle increases, it decreases due to interference effects within the electron cloud. Consequently, heavier atoms with higher electron counts typically have larger scattering factors at low angles. Additionally, this scattering factor influences the intensity of diffraction peaks, directly affecting how clearly atomic positions can be determined from experimental X-ray diffraction data.
Fig.  2: The expert-reviewed version provides essential clarifications about the angular dependence of atomic scattering factors, emphasizes their direct relationship to atomic electron count, and highlights practical implications for interpreting diffraction data, elements often missed in AI-generated explanations.
Question What happens if the path difference between scattered rays is $n\lambda$? Options
(A) They cancel each other
(B) They reinforce each other …
(C) etc.
Expert-Reviewed Supporting Material “When two waves differ in path by $n\lambda$ (where $n$ is an integer), they are perfectly in phase and interfere constructively, resulting in amplitude reinforcement. This is a fundamental principle in wave interference and forms the basis of the Bragg diffraction condition in crystallography.”
Fig.  3: An open-book mode example with expert-reviewed supporting material that clearly explains how a path difference of $n\lambda$ leads to constructive interference, with additional contextual information about its relevance to crystallography. The correct answer is highlighted in green.
2.2 Evaluation Framework: Closed-Book vs. Open-Book
The OPENXRD evaluation framework is designed to assess the reasoning capabilities of language models in crystallography under two distinct conditions: closed-book and open-book modes.
In the closed-book mode, the model receives only the question along with its corresponding multiple-choice options. It must rely entirely on its pre-trained internal knowledge to answer, without access to any additional resources or clarifications. This setting evaluates the model’s intrinsic understanding of crystallographic principles and its ability to perform domain-specific reasoning based solely on its learned representations.
In contrast, the open-book mode supplements the question and answer choices with a brief, domain-relevant textual passage. This supporting material contains essential background information or clarifications pertinent to the question, similar to consulting a short textbook excerpt. The aim is to evaluate the model’s ability to integrate external context with its internal knowledge in order to reason more effectively.
Together, these two evaluation settings provide complementary insights. While the closed-book mode tests baseline domain knowledge and internal consistency, the open-book mode examines the model’s capacity to synthesize and apply external scientific information. This dual-mode structure is particularly valuable for identifying the limits of a model’s independent expertise versus its ability to benefit from curated guidance.
Question + Options
(Closed-Book Input)
Supporting Textual Material
(Open-Book, GPT-4.5 Supporting Knowledge)
Concatenate / Fuse
(Combined Context)
LLM Inference
(Generates Answer)
QA Result
(Final Output)
Fig.  4: Illustration of our open-book QA pipeline for crystallography. In closed-book mode, the model sees only the question (left rotated box). In open-book mode, it also receives domain-specific supporting textual material (right rotated box), which is concatenated and fed to the LLM (center pipeline), producing the final QA result.
2.3 Supporting Textual Material Generation
2.3.1 AI-Generated Supporting Materials (Open-Book, Without Expert Review)
Due to copyright constraints and the impracticality of systematically scanning large volumes of actual textbooks, GPT-4.5 was used to generate short, supportive textual paragraphs for the open-book evaluations.
The generation process followed three key principles. First, each passage summarized fundamental crystallographic concepts relevant to the specific question. Second, care was taken to avoid directly stating the correct answer, ensuring that models must still reason rather than extract solutions verbatim. Third, the length and content were controlled to maintain clarity and focus, typically resulting in brief paragraphs spanning half to one page.
2.3.2 AI-Generated Materials with Expert Review
We engaged three PhD students specializing in crystallography to review and refine the supporting materials. These experts, with four to seven years of research experience in XRD, crystal structure analysis, and materials characterization, were instructed to correct any technical inaccuracies in AI-generated materials, improve explanations with precise domain terminology, ensure comprehensive coverage of relevant concepts, improve clarity while maintaining concision, remove potentially misleading information, and add critical contextual details missing from the original materials. As demonstrated in Figure 2, the expert-reviewed version provides substantial improvements over the initial AI-generated explanations. Furthermore, Figure 3 illustrates a representative example from the multiple-choice question-answering dataset, showcasing the original question, corresponding answer choices, and the AI-generated explanation as revised by an expert.
2.4 Input Context Structuring
We implement a structured input preparation approach via a fusion module. For open-book evaluations, we employ a retrieve-then-read prompt structure where the supporting material precedes the question:
$x=\textsc{Format}(x_{s},x_{q},x_{o})$ (1)
where $x_{s}$ is the supporting textual content, $x_{q}$ represents the question text, and $x_{o}$ indicates the provided answer choices. The Format function organizes these components with clear delimiters and instructional guidance. For closed-book evaluations, the supporting material is omitted ($x_{s}=\emptyset$), while the remaining formatting is preserved.
This structured approach aligns with established retrieval-augmented generation paradigms, wherein external knowledge serves as contextual reference material. Figure 4 demonstrates the architecture of this open-book QA pipeline. Currently, our implementation is exclusively textual; however, this framework can be extended to multimodal data (such as XRD patterns and crystallographic diagrams) as generative vision models become sufficiently advanced.
2.5 Evaluation Metrics
We use several metrics to quantitatively assess model performance across evaluation modes, primarily focusing on accuracy, calculated as:
Accuracy=|{correctly answered questions}||{all $\text{Accuracy}=\frac{|\{\text{correctly answered questions}\}|}{|\{\text{all questions}\}|}\times 100\%$ (2)
For our 217-question benchmark, accuracy is reported both as an aggregated measure across the entire dataset and as a disaggregated measure by specific subtasks. For each model $\bm{M}$ and subtask $t\in\mathcal{T}$, accuracy is computed as:
$\text{Accuracy}_{\bm{M},t}=\frac{|\{q\in Q_{t}:\bm{M}(q)=y_{q}\}|}{|Q_{t}|}\times 100\%$ (3)
where $Q_{t}$ is the set of questions associated with subtask $t$, $y_{q}$ is the correct answer, and $\bm{M}(q)$ denotes the model’s prediction.
To assess the efficacy of open-book augmentation, we measure performance improvement $\Delta$ as:
$\Delta_{\bm{M}}=\text{Accuracy}_{\bm{M}}^{\text{Open-Book}}-\text{Accuracy}_{\bm{M}}^{\text{Closed-Book}}$ (4)
Additionally, we quantify the relative improvement gained through expert-reviewed materials compared to AI-generated supporting materials:
$\Delta_{\bm{M}}^{\text{Expert}}=\text{Accuracy}_{\bm{M}}^{\text{Open-Book, Expert}}-\text{Accuracy}_{\bm{M}}^{\text{Open-Book, AI}}$ (5)
This comprehensive evaluation framework allows us to explore several aspects. We examine the intrinsic crystallographic knowledge inherent in each model, referred to as closed-book performance. We also assess each model’s ability to leverage external knowledge effectively, known as open-book improvement. Additionally, we evaluate the added value provided by expert-reviewed supporting materials, termed expert refinement differential. Furthermore, we analyze model-specific strengths and weaknesses across various crystallographic subtasks.
3 Results
In this section, we present a detailed evaluation of our open-book QA system for crystallography, using a curated set of 217 XRD-related questions. Each question can be answered in two ways: Closed-book mode, where the model relies solely on its internal knowledge, and Open-book mode, where it consults supporting textual material. Our goal is to determine to what extent supporting textual material access bolsters model performance, particularly for advanced domain-specific queries. We conduct our experiments in two phases: first using AI-generated supporting materials provided by GPT-4.5, and then using expert-reviewed versions of these materials to assess whether human domain expertise further enhances performance.
3.1 Setup and Baselines
Models Compared. We evaluate several large language (and vision-language) models. This includes the O-family Models such as O3-mini (O3-mini-2025-01-31) and O1 (O1-2024-12-17), which are designed for scientific reasoning tasks. We also look at various GPT-4 Variants like GPT-4.5-preview (GPT-4.5-preview-2025-02-27), GPT-4-turbo-2024-04-09, GPT-4-preview-0125, and GPT-4-0613. These versions were tested in an interim stage that incorporates "supporting textual material" into GPT-centered language processing. Additionally, we explore pipelines utilizing LLaVA, including LLaVA-v1.6-mistral-7B, LLaVA-v1.5 (7B, 13B), and LLaVA-onevision-QWEN2, which combine a vision encoder (ViT) with an LLM backbone. These models are significant as they use text-only configurations with "supporting textual materials" that include image captions but not the images themselves, within the broader context of this research.
Inference Setup. We employ a zero-shot inference paradigm for all models without parameter updating or fine-tuning on our crystallography evaluation set. For each query, we construct a prompt consisting of the question and enumerated options, formatted consistently across all models. No in-context learning examples or demonstrations are provided, requiring models to rely entirely on their parametric knowledge. All foundation models are evaluated using their published weights and architectures; O-family models have documented parameter-efficient training on scientific corpora, while vision-language models in the LLaVA family were primarily trained on general domain text-image pairs without domain-specific adaptation to crystallography.
Experimental Progression. Our experimental approach follows two distinct phases: In the first phase, we use GPT-4.5 to generate supporting textual materials for each question, carefully prompting it to provide relevant information without revealing the answer. Following this, in the second phase, we engaged three Ph.D. students specializing in crystallography, with 4-7 years of research experience, to review and refine the supporting materials for improved accuracy and pedagogical clarity. This two-phase approach allows us to quantify the value added by human domain expertise in the construction of supporting materials relative to state-of-the-art AI generation.
3.2 Closed-Book Mode Observations
Table 1 presents the closed-book mode accuracy achieved on our 217-question crystallography benchmark, and Figure 5 illustrates the categorized closed-book performance of the models. O3-mini performs extremely well, achieving about 89% accuracy, even surpassing some GPT-4 variants, likely due to targeted domain pre-training. However, smaller LLaVA-based models face challenges with advanced reflection extinctions, often incorrectly identifying which plane indices vanish for BCC or confusing zone-axis notation. Despite these issues, nearly all models correctly answer basic factual questions like the number of crystal systems and who discovered X-ray diffraction.
Table 1: Comprehensive "closed-book mode" accuracy results on the 217-item crystallography QA benchmark, including older GPT-4 variants and newly reported models.
Rank Model (Closed-book mode) Accuracy (%) Correct #
1 GPT-4.5-preview 93.09 202/217
2 O3-mini 88.94 193/217
3 O1 87.56 190/217
4 GPT-4-turbo-preview 83.41 181/217
5 GPT-4 81.11 176/217
6 LLaVA-v1.6-34B 66.80 145/217
7 LLaVA-onevision-QWEN2-7B-si 66.40 144/217
8 LLaVA-onevision-QWEN2-7B-ov-chat 65.90 143/217
9 LLaVA-onevision-QWEN2-7B-ov 65.40 142/217
10 LLaVA-v1.6-mistral-7B 53.00 115/217
11 LLaVA-onevision-QWEN2-0.5B-ov 47.50 103/217
12 LLaVA-v1.5-13B 46.50 101/217
13 LLaVA-onevision-QWEN2-0.5B-si 46.10 100/217
14 LLaVA-v1.5-7B 18.00 39/217
Large ModelsMid-sized ModelsSmall Models88.9%66.2%37.2%Model GroupsClosed-book performance
Fig.  5: Comparison of closed-book performance by model size group. Large Models include GPT-4 variants and O3-mini; Mid-sized Models include LLaVA-34B and QWEN2-7B variants; Small Models include LLaVA-13B, LLaVA-7B, and QWEN2-0.5B variants.
3.3 Open-Book Mode Observations Using AI-Generated Materials (No Expert Review)
Table 2 lists the closed-book mode vs. open-book mode accuracy on our 217-question crystallography benchmark. Here and in subsequent tables, $\Delta$ represents the performance improvement defined in Equation (4), with model subscripts omitted as each row corresponds to a different model. We first evaluate open-book mode accuracy by giving each model relevant supporting textual material generated by GPT-4.5. The largest improvements occur in smaller LLaVA-based models, which gain 4–6% from the supporting textual material clarifying domain specifics. O3-mini sees only a minor jump (+0.46%), possibly because it already "knows" the domain. GPT-4.5-preview slightly decreases in open-book mode (-2.77%), suggesting occasional confusion from extraneous supporting textual material or contradictory references. Figure 6 shows an example where supporting materials have caused confusion. In this case, a model with correct closed-book knowledge has been misled by materials mentioning several related but distinct concepts. Our detailed analysis of the AI-generated supporting materials revealed several limitations: occasional technical inaccuracies in specialized crystallographic terminology, inadequate depth of explanation for particularly complex phenomena, information presented at times being too general to help with specific question nuances, and in some cases, the supporting materials were tangential to the precise knowledge needed for the question. These observations led us to hypothesize that expert-reviewed materials might yield greater performance improvements, particularly for mid-sized models with significant capacity but incomplete domain knowledge.
Question Which scattering mechanism causes damping of the atomic scattering factor at higher angles? Options
(A) Compton scattering
(B) Electron cloud interference
(C) Thermal vibration
(D) Core-level absorption
Model Responses LLaVA-v1.6-34B (Closed-book): B ✓ LLaVA-v1.6-34B (Open-book): C ✗ Supporting Material Excerpt “…The atomic scattering factor decreases with increasing scattering angle due to interference effects within the electron cloud. Additionally, thermal vibrations (Debye-Waller factor) can further dampen the intensity at higher angles, though this is a separate effect from the intrinsic angular dependence…”
Fig.  6: Example of confusion in LLaVA-v1.6-34B due to conflicting supporting material. The model had correct knowledge in closed-book mode but was misled by additional content that mentioned both the correct answer and a distractor as possible factors affecting scattering, albeit in different ways.
Table 2: Comparison of model accuracy in closed-book mode vs. open-book mode with AI-generated supporting materials. $\Delta=$ (Open-book mode) -- (Closed-book mode).
Model Closed-book mode (%) Open-book mode (%) $\Delta$
GPT-4.5-preview 93.09 90.32 -2.77
O3-mini 88.94 89.40 +0.46
O1 87.56 88.02 +0.46
GPT-4-turbo-preview 83.41 82.49 -0.92
GPT-4 81.11 82.03 +0.92
LLaVA-v1.6-34B 66.80 72.81 +4.61
LLaVA-onevision-QWEN2-7B-si 66.40 71.89 +5.53
LLaVA-onevision-QWEN2-7B-ov-chat 65.90 72.35 +6.45
LLaVA-onevision-QWEN2-7B-ov 65.40 71.43 +5.99
LLaVA-v1.6-mistral-7B 53.00 58.53 +5.75
LLaVA-onevision-QWEN2-0.5B-ov 47.50 51.15 +3.68
LLaVA-v1.5-13B 46.50 44.24 -2.30
LLaVA-onevision-QWEN2-0.5B-si 46.10 51.15 +5.07
LLaVA-v1.5-7B 18.00 23.96 +5.99
3.4 Open-Book Mode Observations Using AI-Generated Materials with Expert Review
Building on the findings from the initial open-book evaluations, we introduced expert review to improve the quality and effectiveness of the supporting materials. Three PhD-level crystallography experts were engaged to refine the AI-generated materials. Table 3 compares model accuracy in closed-book and open-book modes using expert-reviewed supporting materials across our 217-question crystallography benchmark. The results show substantially improved performance, particularly among LLaVA-based models.
The largest accuracy gains are observed in LLaVA-v1.6-34B (+11.50%) and LLaVA-v1.6-mistral-7B (+11.10%), highlighting the effectiveness of expert-curated guidance in boosting model reasoning. Several other mid-sized models, such as LLaVA-onevision-QWEN2-7B-ov-chat (+10.10%) and LLaVA-onevision-QWEN2-7B-ov (+8.80%), also demonstrate notable improvements.
Interestingly, GPT-4.5-preview exhibits a slight performance drop (-3.69%) even with expert-reviewed texts. This suggests that highly capable models may already possess sufficient internal knowledge or may be more sensitive to redundant or overly detailed inputs.
Table 3: Comparison of model accuracy in closed-book mode vs. open-book mode with expert-reviewed supporting materials. $\Delta=$ (Open-book mode) -- (Closed-book mode).
Model Closed-book mode (%) Open-book mode (%) $\Delta$
GPT-4.5-preview 93.09 89.40 -3.69
O3-mini 88.94 89.90 +0.96
O1 87.56 88.90 +1.34
GPT-4-turbo-preview 83.41 84.80 +1.39
GPT-4 81.11 82.50 +1.39
LLaVA-v1.6-34B 66.80 78.30 +11.50
LLaVA-onevision-QWEN2-7B-ov-chat 65.90 76.00 +10.10
LLaVA-onevision-QWEN2-7B-si 66.40 75.10 +8.70
LLaVA-onevision-QWEN2-7B-ov 65.40 74.20 +8.80
LLaVA-v1.6-mistral-7B 53.00 64.10 +11.10
LLaVA-onevision-QWEN2-0.5B-ov 47.50 54.40 +6.90
LLaVA-v1.5-13B 46.50 48.40 +1.90
LLaVA-onevision-QWEN2-0.5B-si 46.10 53.90 +7.80
LLaVA-v1.5-7B 18.00 28.60 +10.60
3.5 Comparative Analysis of AI-Generated Supporting Materials: Without vs. With Expert Review
To better understand the impact of expert review, Table 4 presents a direct comparison of performance gains ($\Delta$) between AI-generated supporting materials without expert review and those refined through expert review. Our comparative analysis reveals several significant patterns and Key Findings from Comparative Analysis. A near-universal improvement was observed with expert review, where 13 out of 14 models showed greater performance gains or smaller losses compared to those using AI-generated materials. Specifically, mid-capacity models like LLaVA-v1.6-34B and LLaVA-v1.6-mistral-7B saw the most dramatic improvements, with additional gains of 6.89% and 5.35% respectively. Conversely, models such as LLaVA-v1.5-13B and GPT-4-turbo-preview, which initially showed negative effects with AI-generated materials, experienced positive effects with expert-reviewed ones. However, top models already near ceiling performance, such as GPT-4.5-preview, O3-mini, and O1, saw minimal additional benefits or even slight degradation. Performance differences suggest that expert-reviewed materials successfully addressed the limitations identified in Phase 1, particularly for models with sufficient capacity to utilize improved information but insufficient internal knowledge to solve problems independently. Furthermore, Figure 7 illustrates a categorized performance comparison of these two phases of simulations.
Table 4: Direct comparison of performance gains between AI-generated supporting materials with and without expert review.
Model AI-Generated $\Delta$ (%) Expert-Reviewed $\Delta$ (%) Difference
GPT-4.5-preview -2.77 -3.69 -0.92
O3-mini +0.46 +0.96 +0.50
O1 +0.46 +1.34 +0.88
GPT-4-turbo-preview -0.92 +1.39 +2.31
GPT-4 +0.92 +1.39 +0.47
LLaVA-v1.6-34B +4.61 +11.50 +6.89
LLaVA-onevision-QWEN2-7B-ov-chat +6.45 +10.10 +3.65
LLaVA-onevision-QWEN2-7B-si +5.53 +8.70 +3.17
LLaVA-onevision-QWEN2-7B-ov +5.99 +8.80 +2.81
LLaVA-v1.6-mistral-7B +5.75 +11.10 +5.35
LLaVA-onevision-QWEN2-0.5B-ov +3.68 +6.90 +3.22
LLaVA-v1.5-13B -2.30 +1.90 +4.20
LLaVA-onevision-QWEN2-0.5B-si +5.07 +7.80 +2.73
LLaVA-v1.5-7B +5.99 +10.60 +4.61
Large ModelsMid-sized ModelsSmall Models0.370.37-0.37- 0.37%$5.71$%$3.11$%$0.08$%$9.64$%$6.55$%AI-Generated MaterialsExpert-Reviewed Materials
Fig.  7: Average performance gains from AI-generated vs. expert-reviewed supporting materials across model size categories. Large Models (e.g., GPT-4, GPT-4.5, O1, O3-mini); Mid-sized Models (e.g., LLaVA-34B, QWEN2-7B); Small Models (e.g., LLaVA-13B, LLaVA-7B, QWEN2-0.5B). Note the consistently larger gains achieved with expert-reviewed supporting textual materials, particularly evident in mid-sized models.
3.6 Detailed Subtask-Level Performance with Expert-Reviewed Materials
To gain deeper insights, we examined subtask-level performance with expert-reviewed materials. Table 5 shows how LLaVA-onevision-QWEN2-7B-ov behaves on key XRD subtasks. The rationale for selecting this model is its substantial enhancement from expert-reviewed materials, which offers valuable insights into the limits of performance improvement. For several subtasks such as Crystal Structure, Laue Patterns, Metal Structures, and Powder Diffraction, expert-reviewed materials transformed performance from complete failure (0%) to perfect accuracy (100%). Expert materials particularly improved performance on structural tasks, enhancing Structure Factors by 25% and Coordination Numbers by 40%. However, complex mathematical derivations like Bragg’s Law and Calculation Methods remained challenging even with expert materials, suggesting that some concepts require more than textual explanation. An illustrative example is the question "What is the structure factor $F$ for a base-centered unit cell when $h$ and $k$ are mixed (one even, one odd)?" Correct answer is being $F=f$, while a model consistently selecting $F=2f$, reveals that it has not internalized how lattice symmetry produces systematic absences, an error that echoes its 0% score on Bragg’s Law questions. Additionally, for a few subtasks where the model had strong initial performance, such as Complex Mathematics and Wave Scattering, supporting materials appeared to interfere with the model’s existing knowledge.
Table 5: Selected subtask-level accuracy (%) for LLaVA-onevision-QWEN2-7B-ov, comparing closed-book mode vs. open-book mode with expert-reviewed materials. Some subtasks have fewer questions, so a single item shifts accuracy considerably.
Subtask Closed-book mode Open-book mode Improvement
Large Gains
   Atomic Spacing 66.7 100.0 +33.3
   Coordination Numbers 20.0 60.0 +40.0
   Crystal Structure 0.0 100.0 +100.0
   Laue Patterns 0.0 100.0 +100.0
   Metal Structures 0.0 100.0 +100.0
   Powder Diffraction 0.0 100.0 +100.0
   Structure Factors 60.0 85.0 +25.0
Little / No Improvement
   Bragg’s Law 0.0 0.0 +0.0
   Calculation Methods 0.0 0.0 +0.0
   Diffraction Limitations 0.0 0.0 +0.0
Performance Drops
   Complex Mathematics 100.0 50.0 -50.0
   Wave Scattering 100.0 0.0 -100.0
4 Discussion
Our results show that using domain-specific prompts, or open-book mode, significantly enhances question-answering accuracy in crystallography. This is especially true for smaller or more general models, where the additional context helps fill knowledge gaps. However, the quality and relevance of the supporting textual material are crucial. In some instances, the material generated by GPT-4.5 did not align well with the question’s needs, causing confusion or contradictory information. This underscores the importance of relevance filtering or validation of supporting material before it is used.
Advanced mathematical reasoning remains a challenge. Our analysis indicates that open-book mode rarely resolves complex mathematical problems, such as structure factor calculations or multi-step interference proofs. Future enhancements could include integrating symbolic math modules or domain-specific solvers to better handle these tasks, rather than relying solely on textual references.
The complexity of visual data in crystallography, such as tabulated results, diffraction patterns, and annotated structural diagrams, presents additional challenges. Accurate extraction of this information requires improvements in optical character recognition (OCR), figure parsing, and multi-modal alignment. Developing tailored visual backbones or domain-tuned OCR engines will be essential for applying our methods to real-world scenarios, like analyzing lab notebooks or older textbook scans.
In parallel, a notable tension exists between generalist and specialist models. While larger LLMs like GPT-4 and GPT-4.5 have broad knowledge, they often miss the nuances of specific domains that specialized models capture more effectively. A hybrid approach, where a specialist model generates domain-specific references for a generalist model, could offer a powerful solution. Additionally, ethical and copyright considerations must be managed carefully, especially as we look to expand these systems to other technical fields with extensive proprietary literature.
Our research identifies two primary methods for enhancing model performance with domain expertise: directly embedding knowledge into large models, such as GPT-4.5 and O3-mini, and providing models with access to expert-reviewed references during inference. This latter strategy proves especially useful for mid-capacity models, enabling them to nearly match the performance of their larger counterparts by leveraging a "knowledge bridge."
For example, a medium-sized model like LLaVA-v1.6-34B, when paired with expert materials, delivers performance close to that of GPT-4 but with far fewer resources required. The integration of expert-reviewed domain knowledge is not only beneficial in textual analysis but also in multi-modal contexts such as medical imaging or geospatial analysis, where expert-reviewed context can similarly boost specialized question-answering performance and unlock new application areas.
However, the addition of external information is not always beneficial. For instance, GPT-4.5-preview shows a decline in performance when supplemented with external materials, possibly due to conflicts with its pre-trained internal knowledge. This suggests that more information is not necessarily better, particularly for top-tier models that have been extensively trained.
Future research should focus on automated methods to determine when expert review is most beneficial, multi-stage approaches for targeted expert review, and the potential for fine-tuning models to better utilize external references. Extending these methodologies to other specialized domains could further validate the effectiveness of this approach.
Despite these advancements, certain limitations remain. Complex mathematical concepts and advanced tasks like reflection or symmetry analysis still pose challenges, as models do not always fully grasp these concepts through text-based materials alone. Moreover, there are instances where supporting materials can detract from a model’s performance if they do not align perfectly with the model’s pre-existing knowledge base.
Given these limitations, the reliance of open-book question answering on human-curated content may be difficult to sustain across all fields, and the uneven gains observed across model sizes underscore the need for further work on how best to select, refine, and integrate external knowledge.
5 Conclusions
We present OPENXRD, an open-book question-answering framework for crystallography that shows how AI-generated supporting materials can improve model performance on specialized scientific questions. Through evaluation on 217 expert-level XRD questions, we find that mid-capacity models (7B-34B parameters) benefit most from external knowledge, gaining up to +11.5 percentage points with expert-reviewed materials compared to +6% with AI-generated content alone. Our main finding is an inverted-U relationship: the largest models show minimal improvement or slight degradation with external materials, while the smallest models have limited capacity to utilize additional information effectively. Mid-sized models occupy a "sweet spot" where they have sufficient processing capability but incomplete domain knowledge. The practical value lies in demonstrating that carefully curated supporting materials can help smaller models approach larger model performance on specialized tasks. This suggests cost-effective deployment strategies where mid-sized models with expert-reviewed materials can substitute for much larger, more expensive models in domain-specific applications. We release our dataset and code to enable further research on how different AI systems utilize external knowledge in specialized domains. While our work focuses on crystallography, the methodology could extend to other technical fields requiring expert-level question-answering.
Author contributions
A.V. and A.S. contributed equally to this work as co-first authors. A.V.: Conceptualization, Methodology, Software, Formal analysis, Investigation, Data curation, Writing – original draft, Visualization. A.S.: Conceptualization, Methodology, Data curation, Investigation, Validation, Writing – review & editing. A.S. designed the crystallography question set and led the expert review process for GPT-generated questions. Z.Z., Y.X., and G.H.: Methodology, Software, Investigation, Validation, Writing – review & editing. Z.Z., Y.X., and G.H. provided substantial contributions to model development, experimental design improvements, and participated in the expert review and refinement of GPT-generated questions to ensure scientific accuracy. C.X. and N.A.: Conceptualization, Supervision, Project administration, Resources, Writing – review & editing, Funding acquisition. C.X. and N.A. jointly led and supervised the project, providing strategic direction and ensuring research quality throughout all phases of the work.
Conflicts of interest
The authors declare that they have no competing financial interests or personal relationships that could have appeared to influence the work reported in this document.
Data availability
The complete dataset of 217 expert-curated crystallography questions, supporting materials (both AI-generated and expert-reviewed versions), evaluation scripts, and model implementation code are publicly available at https://github.com/niaz60/OpenXRD. Additional project information and resources can be found at https://niaz60.github.io/OpenXRD/. The repository includes detailed documentation for reproducing all experimental results presented in this study.
Acknowledgments
We gratefully acknowledge financial support from multiple sources. A.V., G.H., C.X., and N.A. received support from the National Nuclear Security Administration (NNSA) under grant NA0004078. A.S., Z.Z., Y.X., C.X., and N.A. were supported by the National Science Foundation (NSF) under grant 2202124. Additional funding for N.A. was provided by the Department of Energy (DOE) under award DE-SC0020340. We thank the crystallography experts who participated in the review and validation of our question dataset. We also acknowledge the computational resources and infrastructure that enabled the extensive model evaluations presented in this work.